L2=75 mm L3=170 mm cut stroke 45 mm OXO 0,05 workpiece

How many degrees of freedom does this machine have?
Answer #1

The degree of freedom is

n = 3(L-1) - 2J - h

There are total 6 links. L = 6

There are total 5 joints. j = 5

The cutter will be having point contact with workpiece initially.

So the number of higher pairs will be 1. h = 1.

Now,

n = 3(L-1) - 2J - h

n = 3(6-1) - 2(5) - 1

n = 4.

There are 4 degrees of freedom.